# Unreal Robotics Lab: A High-Fidelity Robotics Simulator with Advanced Physics and Rendering

### v2

@@ I INTRODUCTION
Simulation plays a crucial role in robotics research, enabling safe, efficient, and cost-effective development, testing, and validation of robotic systems before real-world deployment. High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality. However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [szot2021habitat]. This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [unrealengine] with the advanced physics simulation of MuJoCo [todorov2012mujoco] to create photorealistic environments for robotics applications. By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects — such as smoke, fire, and water waves — which are crucial to training robots to handle diverse and unpredictable conditions. These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing general-purpose robotics simulators. As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
@@
We present a novel, open-source simulation framework — the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation. Unlike existing simulators focused on specific domains (e.g. autonomous driving or aerial robotics) or embodied AI for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more. While powerful general-purpose simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, our framework prioritizes physical fidelity by utilizing the MuJoCo physics engine, which is highly validated for its accuracy and stability in contact-rich control tasks. Additionally, we incorporate CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation. Our key contributions are:
  • High-Fidelity Physics with Photorealistic Rendering: Synchronization of MuJoCo’s physics with Unreal Engine’s advanced rendering (Lumen, Nanite, Niagara) enables realistic perception-based training and benchmarking.
  • Generalized Benchmarking for Diverse Conditions: A structured benchmarking system evaluates robotic perception and control under rare conditions (e.g., smoke, fog, dynamic lighting, dynamic movement).
  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with Robot Operating System (ROS) compatibility.
  • Open-Source and Extensible: The system will be released as open-source, supporting modular extensions for new controllers, planners, and robotic platforms.
@@ II RELATED WORK
Robotic simulation has advanced significantly in recent years, with various frameworks addressing key challenges such as physics accuracy, photorealistic rendering, and real-time performance. This section reviews related work in three main areas: robotics-focused simulation platforms, comparisons of physics engines, and benchmarking SLAM and navigation under adverse visual conditions.
Several other game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry. Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [unrealengine] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change. Furthermore, Nanite’s automatic level of detail (LOD) [unrealengine] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance. This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical. Beyond rendering, Unreal Engine offers a flexible and extensible ecosystem, supporting third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment, making it highly adaptable for a variety of robotic simulation tasks. These features ensure that simulations can achieve both high physical accuracy and photorealistic perception, making them an ideal choice for vision-based robotics research.
@@ II-A Robotics-Focused Simulation Platforms
Robotic simulation has made considerable progress, with numerous frameworks addressing challenges in physics accuracy, rendering quality, and real-time performance. This section reviews related work in three primary areas: robotics simulation platforms, comparisons of physics engines, and benchmarking of SLAM and navigation under adverse visual conditions. Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency. Gazebo [koenig2004design] is a widely-used open-source simulator that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], (Py)Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering capabilities lag behind modern game-engine-based simulators. MuJoCo [todorov2012mujoco] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering. GPU-accelerated platforms such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] enable large-scale reinforcement learning, prioritizing speed over advanced rendering. Isaac Sim [NVIDIA_Isaac_Sim], built on NVIDIA Omniverse, provides extensible robotics simulation with high-fidelity physics via PhysX [nvidia_physx]—which excels in angular stability and scalability but may lag in accuracy for contact dynamics and linear stability compared to MuJoCo. Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments suitable for embodied AI research. CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for photorealistic urban and aerial simulations, with a focus on autonomous driving and drone applications. CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding. However, both simulators are largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios. RoboSuite [robosuite2020] and RoboCasa [robocasa2024] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity. Although most simulators prioritize either physics accuracy or visual realism, few achieve both. Our approach bridges this gap by integrating MuJoCo (precise physics) with Unreal Engine (state-of-the-art rendering), creating a high-fidelity simulation pipeline for diverse robotic applications.
@@ II-B Physics Engines for Robotics Simulation
Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions. Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [erez2015simulation]. MuJoCo [todorov2012mujoco] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion. However, it lacks built-in high-fidelity rendering. ODE (Open Dynamics Engine) [koenig2004design], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning. Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
Unreal Engine’s Chaos and Havok physics solvers [havok_physics] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations. However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics. NVIDIA PhysX [nvidia_physx] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control. Our approach combines MuJoCo’s precise physics, which are validated for their accuracy and stability [erez2015simulation], with Unreal Engine’s advanced rendering. This integration achieves both photorealism and accurate simulation for robotic applications. It also provides a broad feature set by inheriting the built-in capabilities of both Unreal Engine and MuJoCo.
@@ II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting. Envodat [nwankwo2024envodat] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity. Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings. Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim). While our work shares a similar goal with recent simulators such as SPEAR [simulator_intelspear] simulator in combining Unreal Engine with MuJoCo, our frameworks are built on fundamentally different design philosophies. SPEAR is primarily structured as a Python-based OpenAI Gym environment, making it well-suited for task-specific, algorithmic research in embodied AI. In contrast, our framework is designed for systems-level validation through its native ROS integration, allowing researchers to test a complete robotics software stack as it would be deployed in the real world. This focus on flexible, real-world system replication is further reflected in our integrated workflow. Our framework allows users to build and simulate scenes directly within the Unreal Editor, automatically handling all asset processing. This streamlined approach eliminates the need for separate scripts and manual data handling for importing and exporting assets, which can be required in other simulators. Ultimately, this design makes URL a generalized platform where users can freely import any MuJoCo-compatible robot and construct custom tasks, rather than focusing on a curated set of pre-built scenes and benchmarks. Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios. Beyond rendering, UE’s mature ecosystem accelerates robotics research. We utilize its built-in NavMesh and Behavior Tree tools to script complex, dynamic scenarios with moving NPCs and vehicles, creating environments that are not just visually realistic but also behaviorally rich. This is critical for testing planner robustness in non-static worlds, a feature often lacking in physics-focused simulators.
III Simulator setup
III-A Overview
Our simulator is implemented as a UE plugin centered around a custom SimManager Actor, with the overall architecture shown in Fig. 2. This Actor manages several core sub-systems: a SceneManager for processing assets and constructing the MuJoCo scene, a PhysicsManager to synchronize MuJoCo physics with UE rendering, and Benchmark and Replay systems for performance evaluation and data logging. The pipeline begins with the SceneManager scanning for actors with a UCustomPhysicsComponent, processing their properties to construct a MuJoCo Specification (MjSpec), and compiling the final physics model. The system supports standard robots defined via MuJoCo XML with corresponding UE assets, including sensors such as cameras, IMUs, and force sensors. Communication leverages a modified ROS plugin [mania19scenarios] for compatibility with existing robotics frameworks, allowing the simulator to act as a high-fidelity, real-world surrogate for testing robotics behaviors and perception. To complement the physics and data pipeline, our framework uses the larger set of features of Unreal Engine to build rich and dynamic worlds. This includes high-fidelity rendering for perception research (Lumen, Nanite), rapid world-building with automated NavMesh generation for pathfinding, and complex scenario creation using Behavior Trees for intelligent agents and the Niagara system for adverse weather and effects.
III-B Scene Building and Asset Processing
UE actors are configured for MuJoCo using a UCustomPhysicsComponent, which contains an intermediate PhysicsObject to synchronize geometry and state between the two engines. The settings of this component dictate how the object is processed for simulation. Its geometric complexity setting (Complex or Simple) determines how meshes are converted into convex hulls for MuJoCo’s collision system, using either CoACD [wei2022coacd] for decomposition or direct compiler approximation. The component also controls the object’s mobility (static or dynamic) via an override or by defaulting to the actor’s native UE setting.
@@ III-B Scene Building and Asset Processing
The SceneManger then iterates over all actors with the UCustomPhysicsComponent and processes the assets with their containing parameters. Meshes that are processed with CoACD are saved in the project’s SavedDir for caching. UE landscapes are converted to MuJoCo height fields using multithreaded raycasts within a user-defined bounding box at a set resolution, as shown in Fig. 1. This targeted approach avoids processing entire large-scale terrains while still capturing fine details such as small clutter, preserving key environmental features for the physics simulation.
Figure 2: Unreal - MuJoCo Simulator system diagram
MjSpec is then used to generate the final MuJoCo Model and XML file. Robotic assets, packaged as UE pawns with their own XML, are automatically detected and added into the MjSpec. Sensors are defined as sites in the robot’s XML, and we employ a hybrid sensing approach: non-visual data (IMUs, force sensors) is sourced directly from MuJoCo, while visual data (cameras, LiDAR) is rendered in UE for photorealism. The exported XML and assets can also be used to run the scene as a standalone MuJoCo simulation, independent of UE.
III-C Physics Manager
The physics simulation operates asynchronously, running within a separate execution thread to maximize performance. Specifically, the MuJoCo physics state is advanced by calling mj_step within a dedicated thread that maintains a configurable physics timestep (e.g., 1000 Hz) by yielding control between steps. Meanwhile, the Unreal Engine rendering thread operates independently at its own rate, typically 60 FPS or higher, depending on scene complexity. Transformation updates are managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo. To ensure data integrity, locking mechanisms are used when the rendering thread accesses the latest physics state, a design that mimics the behavior of MuJoCo’s native simulate GUI and decouples rendering performance from physics fidelity. External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment. UE physics can still be used for objects not managed by MuJoCo.
III-D Communication
Our framework uses a modified ROS plugin [mania19scenarios] that automates the communication layer. This plugin automatically publishes all sensor data from both MuJoCo and UE to their corresponding ROS topics, the names of which are derived directly from the sensor definitions in the robot’s XML file. Similarly, for robot control, the plugin dynamically generates subscribers for any defined actuators or joints. This allows standard robotics controllers to send actuator or joint commands directly to MuJoCo.
III-E Benchmark System
The benchmarking system automates metric recording and management through a structured, extensible framework built upon a MetricManager and a BaseMetric class. The BaseMetric class provides a fundamental interface for all metrics, handling time-series data storage and defining a consistent lifecycle with methods for initialization, reset, and CSV export. New metrics can be created by deriving from this class to implement domain-specific logic while inheriting the core data-handling functionality. The MetricManager automatically discovers and registers all metric classes by scanning project definitions, allowing users to select and configure them for each experiment via a UI. By default, the system tracks core metrics such as DistanceToGoal (distance to target), Collisions (contact events), TimeToGoal (task completion time), and GlobalPose (absolute pose). All metrics are recorded at discrete intervals and saved as CSV files in designated experiment folders, with support for configurable grace periods and timeouts to ensure reproducible results.
III-F Replay System
Our system has an automated recording and replay mechanism for sensor and body pose data, allowing efficient data collection and experimentation. It uses custom structs to track only essential information, pose data, sensor readings, and timestamps, storing them in a queue for minimal overhead and precise sequence reproduction. Data streams are managed through separate JSON files that can be dynamically saved and loaded at runtime. By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines. This allows researchers to collect data once and replay them in visually modified environments, supporting domain adaptation, robustness testing, and dataset enhancement for machine learning and benchmarking.
III-G Example Deployment of Robots, Controllers, and Visual Planners

### v3

@@ I INTRODUCTION
Simulation is essential for robotics research, enabling safe, efficient development, testing, and validation of robotic systems before real-world deployment. High-fidelity environments must simultaneously deliver realistic perception and accurate physics interactions—an open challenge due to competing computational demands and the complexity of real-world dynamics [szot2021habitat].
We introduce the Unreal Robotics Lab (URL), a novel open-source framework that tightly integrates Unreal Engine’s state-of-the-art rendering pipeline (Lumen global illumination, Nanite virtualized geometry, Niagara particles) [unrealengine] with MuJoCo’s validated high-precision physics engine [todorov2012mujoco]. By embedding MuJoCo natively inside the Unreal ecosystem via shared-memory execution, URL achieves deterministic physics while exposing Unreal’s full authoring, environmental-effect, and tools directly to roboticists.
A key capability is the effortless generation of complex dynamic phenomena — smoke, fire, water waves, dynamic lighting, and moving agents—that are rare or hazardous to capture in real datasets. These effects enable controlled benchmarking of vision-based policies under out-of-distribution conditions and the creation of large-scale training data for perception and navigation.
@@
Unlike domain-specific simulators (e.g., autonomous driving or aerial robotics) or platforms focused on limited embodied-AI scenarios, URL is designed for general-purpose robotic applications, supporting quadrupeds, mobile manipulators, UAVs, and arms. While simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, URL prioritises physical fidelity by retaining MuJoCo’s superior accuracy and stability in contact-rich tasks and incorporates CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation.
Our key contributions are:
  • Native “Unreal-First” Architecture: We introduce a shared-memory execution model that embeds the MuJoCo engine directly within the Unreal Engine environment. This eliminates the latency and synchronization overhead of traditional client-server models.
  • High-Fidelity Physics-to-Visual Synchronization: By mirroring the MuJoCo schema within the UE Actor-Component hierarchy, we enable seamless synchronization between MuJoCo’s stable physics and UE’s photorealistic rendering pipeline (Lumen, Nanite, Niagara).
  • Automated Environmental Processing and Geometry Conversion: The framework provides a robust pipeline for bridging complex UE environmental assets to physics-ready formats, utilizing multi-threaded terrain sampling and automated CoACD-based conversion for arbitrary static meshes.
  • Comprehensive ROS-Compatible Middleware: A unified interface for sensors and actuators that supports diverse robotic platforms with native support for ROS and ZeroMQ communication. This enables the “plug-and-play” deployment of modern planners (e.g., ViNT, NoMAD) and SLAM frameworks (e.g., ORB-SLAM3) directly within the simulation.
  • Generalized Benchmarking for Diverse Conditions: We provide a structured framework for evaluating robotic systems under rare or extreme environmental conditions, complemented by a deterministic replay system that allows for robust domain randomization testing across varied visual environments.
@@ II RELATED WORK
Robotic simulation has advanced significantly in recent years, addressing key challenges in physics accuracy, photorealistic rendering, and real-time performance. This section reviews related work in three main areas: robotics-focused simulation platforms, physics engine comparisons, and benchmarking of SLAM and navigation under adverse visual conditions.
Game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], yet Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry [unrealengine]. Lumen enables real-time global illumination and reflections without precomputed lighting—ideal for dynamic environments—while Nanite efficiently renders highly detailed assets with millions of triangles. Combined with its extensible ecosystem (third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment), Unreal Engine is particularly well-suited for vision-based robotics research.
@@ II-A Robotics-Focused Simulation Platforms
Several simulators target robotic perception, control, and reinforcement learning [ellis2022navigation, shanks2025dreamernav], each with different trade-offs in realism and efficiency. Gazebo [koenig2004design] is a widely-used open-source platform that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering lags behind modern game-engine solutions. MuJoCo [todorov2012mujoco] excels in accurate, efficient physics for control tasks but lacks built-in photorealistic rendering. GPU-accelerated frameworks such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] prioritise speed for large-scale reinforcement learning, while Isaac Sim [NVIDIA_Isaac_Sim] (built on NVIDIA Omniverse) uses PhysX [nvidia_physx] for scalability at the expense of contact-dynamics accuracy compared with MuJoCo. Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments for embodied AI, and domain-specific simulators such as CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for urban/aerial scenarios. Manipulation-focused suites RoboSuite [robosuite2020] and RoboCasa [robocasa2024] rely on MuJoCo physics but offer limited environmental complexity.
Although most platforms emphasise either physics fidelity or visual realism, few achieve both. URL bridges this gap by natively integrating MuJoCo’s precise physics with Unreal Engine’s state-of-the-art rendering in an editor-first, ROS-native framework designed for general-purpose robotics (quadrupeds, mobile manipulators, UAVs, and arms).
@@ II-B Physics Engines for Robotics Simulation
Physics engines are critical for realistic contact, soft-body, and fluid modelling. MuJoCo [todorov2012mujoco] is widely favored for its for soft-contact accuracy and stability in complex robotic tasks [erez2015simulation]. ODE (used in Gazebo) suffers from stability issues in high-DOF scenarios and requires extensive tuning. Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation but lags MuJoCo in precision. NVIDIA PhysX [nvidia_physx] offers GPU acceleration suitable for fast RL yet omits Coriolis forces and exhibits lower linear-stability accuracy. Unreal Engine’s Chaos and Havok solvers [havok_physics] excel at real-time destruction and particle effects but are not optimised for high-precision joint dynamics or soft-contact robotics.
The continued adoption of MuJoCo’s physics is further evidenced by the fact that next-generation GPU-accelerated simulators such as Genesis [Genesis] and Newton [newton_physics] have adopted MJWarp [mujoco_warp] (a high-performance GPU backend directly derived from MuJoCo) as their core physics engine. This widespread use in the latest state-of-the-art platforms confirms that MuJoCo’s contact modelling, accuracy, and stability remain unmatched—even as the field shifts toward massive parallel GPU simulation.
URL therefore retains native MuJoCo as the sole physics backend while inheriting Unreal Engine’s advanced rendering, particle systems, and tools—delivering both photorealism and validated physical fidelity without compromise.
@@ II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
Evaluating SLAM and navigation algorithms under adverse conditions (smoke, fog, dynamic lighting, or moving agents) remains difficult, as most public datasets and simulators capture only structured, well-lit scenes [EmbleyRiches2025SAGESA]. Envodat [nwankwo2024envodat] provides one challenging real-world dataset, yet scale and diversity are still limited. URL addresses this limitation by synthetically generating controllable rare-event scenarios via Niagara particles and fluids, together with dynamic NPCs and vehicles scripted through Unreal’s NavMesh and Behavior Trees.
Current simulators typically prioritise either high-precision physics (e.g., MuJoCo-based) or photorealistic rendering (e.g., CARLA, AirSim). While our work shares the high-level goal of SPEAR [simulator_intelspear] in combining Unreal rendering with MuJoCo physics, the design philosophies differ fundamentally. SPEAR is structured as a Python-based OpenAI Gym environment for task-specific embodied-AI research. In contrast, URL is built for systems-level validation: it offers native ROS integration, direct scene authoring inside the Unreal Editor (with automatic asset processing and CoACD convex decomposition), and seamless import of any MuJoCo-compatible robot. This enables full-stack testing of complete robotic software pipelines exactly as they would be deployed in the real world, together with repeatable benchmarking of perception and planning robustness in behaviorally rich, non-static environments.
III Simulator Architecture
III-A System Overview
We propose an “Unreal Engine-first” simulation framework (Fig. 1) that embeds the MuJoCo physics engine entirely within the Unreal Engine (UE) ecosystem. By discarding traditional client-server communication models in favor of in-process, shared-memory execution, the architecture ensures absolute physical determinism. This design enables the physics simulation to natively leverage UE’s photorealistic rendering pipeline (Lumen, Nanite), environmental effects (Niagara), and AI navigation (NavMesh) without the latency or synchronization overhead inherent in distributed architectures.
Figure 1: Unreal - MuJoCo Simulator system diagram.
III-B World Authoring and Asset Instantiation
A core principle of this framework is the comprehensive mirroring of the MuJoCo schema within UE’s Actor-Component hierarchy. Rather than providing an abstracted wrapper, the system exposes nearly the entire MuJoCo specification—including global defaults, solver options, and runtime configurations—as native, strongly-typed properties directly within the UE interface. The framework caters to both legacy MuJoCo users and native Unreal Engine development through a tiered ingestion pipeline.
Users can integrate legacy models by dragging standard MuJoCo XML files directly into the UE Content Browser. This triggers a dedicated factory that parses the MJCF specification to automatically generate a structurally equivalent UE Blueprint. Articulated robots and environmental objects are instantiated as AMjArticulation Actors, which manage a recursive tree of localized components such as UMjBody, UMjGeom, UMjJoint, and UMjTendon. This process automatically imports and converts all referenced meshes and primitives into native UE assets, creating a persistent representation that benefits from UE’s advanced material and rendering systems.
For native authoring, users can instantiate an empty AMjArticulation and procedurally build the hierarchy through the standard UE “Add” interface. The architecture abstracts away from ID-based indexing and manual pointer arithmetic; users can access live physics data directly through the articulation’s component references without managing raw mjModel or mjData addresses. Scene construction is finalized by placing these Blueprints within a level containing an AMjManager Actor. Upon initialization, the manager performs a scene-wide scan to aggregate all active Actors into a contiguous specification. To optimize performance, the framework utilizes asynchronous asset serialization and caching for mesh generation, ensuring complex geometries are processed once. The system maintains bidirectional compatibility, allowing users to export UE-authored levels back to standardized MuJoCo formats.
III-C Environmental Sampling and Geometry Processing
To bridge the gap between UE’s diverse environmental assets and MuJoCo’s collision requirements, the framework provides automated conversion utilities. The AMjHeightfieldActor facilitates terrain interaction by executing multi-threaded, bounded raycasts to sample the environment. This actor is not restricted to UE Landscapes; it can sample any geometry not already managed by the MuJoCo solver, specifically any Actor lacking an AMjArticulation or a UMjQuickConvertComponent. This allows for the generation of optimized MuJoCo heightfields from complex, non-articulated environmental geometry while minimizing memory overhead (Fig. 2). For discrete objects, the UMjQuickConvertComponent enables the instant conversion of arbitrary UE static meshes into MuJoCo-compatible collision bodies using CoACD [wei2022coacd] for automatic convex hull decomposition.
@@ III-C Environmental Sampling and Geometry Processing
III-D Unified Interaction and Bi-Directional Physics
To maintain strict determinism, the MuJoCo simulation executes on a dedicated asynchronous thread. The solver steps at the specific timestep defined within the MuJoCo model, which is exposed and configurable via the AMjManager. Interaction is unified across the framework’s API; every MuJoCo property, function, and sensor is accessible via both C++ and UE Blueprints. This includes a comprehensive sensor interface where all MuJoCo-native sensors are exposed as UE components, providing researchers with Blueprint nodes to access real-time data streams. Furthermore, the framework includes a native graphical UI (MjSimulateWidget) that replicates the functionality of MuJoCo’s standalone Simulate, allowing for real-time visualization, state monitoring, and interactive control within the UE viewport. While the framework handles pointer management, users retain the ability to interact with the raw MuJoCo structs if required for custom solver logic. This flexibility is complemented by a robust one-way kinematic coupling (mocap) which allows UE’s native Chaos physics and AI systems to influence the MuJoCo solver in real-time. For example, an NPC governed by the UE NavMesh can be assigned a UMjQuickConvertComponent with its “Unreal Driven” property enabled, transforming it into a mocap body that moves in synchronization with UE’s physics.
III-E Communication, Metrics, and Replay
The framework supports multiple communication modalities, allowing users to choose between a modified ROS plugin [mania19scenarios] or a high-performance ZeroMQ (ZMQ) bridge. These plugins automate the mapping of MuJoCo and UE sensors to their respective topics or sockets based on the robot’s configuration. For empirical evaluation, a MetricManager provides a structured framework for recording time-series data directly to CSV. Furthermore, a high-fidelity Replay System caches joint kinematics and sensor streams, enabling researchers to record physical trajectories and deterministically replay them across iteratively augmented visual environments for robust domain randomization testing.
III-F Example Deployment of Robots, Controllers, and Visual Planners